OpenStudy (dls):

If the KE of the particle is increased 16 times the percentage change in deBroglie wavelength of the particle is?

OpenStudy (dls):

\[\LARGE \lambda=\frac{h}{mv}=\frac{h}{\sqrt{2m KE}}=\frac{h}{\sqrt{2m (KE+16KE)}}\]

OpenStudy (anonymous):

!? why did you add? :O

OpenStudy (dls):

increased 16 times o.o

OpenStudy (dls):

first it was x.. after increasing 16 times.. 16x+x=17x :"D

OpenStudy (anonymous):

its said increased 16 times.. do you know the meaning of "times" ?!

OpenStudy (anonymous):

5 times 2 is not 5+2.... :P

OpenStudy (dls):

so .. \[\Huge \lambda=\frac{h}{4 \sqrt{2 KE m}}\]

OpenStudy (dls):

so it becomes 1/4 times.. % decrease: 25% so new one is 75% i got this answer earlier too but its wrong..answer isn't that.. @Mashy

OpenStudy (dls):

@Mashy

OpenStudy (anonymous):

it decreases 75%.. the new one is 1/4times.. means the new is one 25 %

OpenStudy (dls):

yes i got the same 25% wahi but ans is 60%!!

OpenStudy (dls):

(verified ans)

OpenStudy (dls):

that is what confused me :(

OpenStudy (anonymous):

thats preposterous

OpenStudy (dls):

A resistance of 20 ohms is connected to a source of ac rated 110V,50Hz. Find the time taken by the current to change from its max to rms value?? @Mashy

OpenStudy (anonymous):

just use the the current equation..

OpenStudy (dls):

how??

OpenStudy (anonymous):

whats the instantaneous current equation!?

OpenStudy (dls):

\[\LARGE i=i_0 \sin (\omega t+\phi)\]

OpenStudy (anonymous):

its pure resistive.. you can calculate i_0 and omega.. right?!